If you place a 80-foot ladder against the top of a 64-foot building, how many feet will the bottom of the ladder be from the bottom of the building?
Found 2 solutions by rothauserc, MathTherapy:
Answer by rothauserc(4718)

This problem can be modeled with a right triangle whose hypotenuse is 80 feet, one side is 64 feet, and the distance from the bottom of the building is side x
:
Apply the Pythagorean Theorem for a right triangle
:
64^2 +x^2 = 80^2
:
x^2 = 80^2 -64^2 = 6400 - 4096 = 2304
:
take square root of both sides of =
:
x = 48
:
*******************************************************************
The bottom of the ladder is 48 feet from the bottom of the building
*******************************************************************
:

Answer by MathTherapy(10552)


If you place a 80-foot ladder against the top of a 64-foot building, how many feet will the bottom of the ladder be from the bottom of the building?
This represents one of the PYTHAGOREAN TRIPLES, namely: (3-4-5), but TIMES 16.
This results in sides: (48-64-80), so missing side is?